OpenStudy (anonymous):

Of the 750 participants in a professional meeting, 450 are female and 1 2 of the female and 1 4 of the male participants are less than thirty years old. If one of the participants will be randomly selected to receive a prize, what is the probability that the person selected will be less than thirty years old?

OpenStudy (ganpat):

750 participant, 450 female , so 750 - 450 = 300 male.. 1/2 of female = 450 / 2 = 225 female are below 30 1/4 of male = 300 / 4 = 75 male are below 30 so probability will be 1 / (225 + 75) = 1/ 300